# Understanding Nonlinear Collaboration between Human and AI Agents: A Co-design Framework for Creative Design

### 14 Jan 2024 v1

@@ 5.4. Lessons We Learnt from Study II
What altered people’s preferences towards OptiMuse and Copilot? To answer this problem, we summarized people’s different expectations of AI agents in human-AI co-design and their preferences. Specifically, we try to address two questions inspired by Li et al. (li2023ai, ): What tasks are anticipated from AI agents, and what roles are expected of AI agents?
@@ 5.4.1. What tasks are anticipated from AI agents?
We adjusted the sequence of these two questions because tasks will be frequently discussed in formulating the definition of AI agents’ roles (Tab. 3 and Tab. 4). The tasks are as follows: Requirement Understanding: comprehend users’ design requirements conveyed through textual input; Requirement Analysis: engage in a dialogue with users to establish consensus regarding design objectives or operations. In other words, provide textual feedback towards users’ words; Solution Exploration: present alternative design options in either textual or visual formats; Review: evaluate designs using design knowledge, making comparisons with existing solutions; Execution: transform operational requirements into visual outputs.
@@ 5.4.2. What roles are expected of AI agents?
Based on the interviews, our users expressed varying preferences and requirements regarding OptiMuse and Copilot. Specifically, they held different perceptions of the roles AI should play in graphic design. We categorized these roles as expert, colleague, optimizer, and executor, noting a diminished emphasis on creativity (Tab. 3).
@@
  • Executor: Do just as I tell you. Among the 3/12 participants, the prevalent choice was for AI to work only as an executor. An executor AI is expected to faithfully carry out users’ commands. This specific type of user has a clear understanding of their goals and the possible operations required to achieve them, although they remain uncertain about the outcome until they actually put a hand to it. As a result, they depend on AI to reduce the manual effort involved in turning operations into visible results. They prefer to avoid requirement analysis, finding it cumbersome to explain operations to AI. Furthermore, they have no inclination to make selective choices because they wish to avoid the additional time and effort required for comparisons and selections among choices.
@@ 6. Discussion
6.1. The Nonlinear Design in Human-AI Co-design Tools
@@
In popular tools like Copilot (copilot, ) and Midjourney (midjourney, ), AI agents necessitate precise instructions and do not facilitate the remixing of various options (Fig. 3). They prioritize addressing the constraints inherent in the human-human co-design process (DR4, DR5). As described in Kulkarni et al.’s work (kulkarni2023word, ), designers often refine their prompts multiple times and regenerate images until they achieve relatively satisfactory results. Following this, they may choose to adopt either a portion or the entire outcome, either through manual editing or by blending prompts to generate new designs. Such turn-based linear workflow hinders the natural nonlinear navigation among creative design stages (hatchuel_new_2003, ; pugh_total_1991, ; designcouncil, ; gero1990design, ; gero2004situated, ; howard_describing_2008, ), which involves drawing inspiration from various parallel iterations and potentially modifying initial ideas along the way (Girotto2016collective, ). As a result, existing AI-assisted tools fail to consider the design requirements (Table.  1) because of their requirements for precise commands (DR1-2) and restricted operations on outcomes (DR3).
@@
Existing frameworks (negrete2014apprentice, ; kantosalo2016modes, ; guzdial2019interaction, ) on human-AI collaboration have shifted attention from AI as mere tools to more collaborative and creative agents. However, they simply describe the co-design process as “turns” or “iteration”. We bridge this gap by formulating a nonlinear human-AI co-design framework featuring flexible communication and selective choices. Following the nonlinear framework, OptiMuse enables parallel iteration (Girotto2016collective, ) and remixing (tan2022conflict, ; Cheliotis2014remix, ; Oehlberg2015remix, ; yue2019improving, ). Similar to Rezwana’s finding in Penpal (rezwana2022understanding, ), we observe enhanced user engagement in OptiMuse, where AI was perceived as more collaborative than a mere tool. In summary, our framework serves as a compelling example of integrating theories (creative design process, human-AI framework) and practices (collective creativity strategies, Study I), which offers practical guidelines for the development of future human-AI co-creative systems.
6.2. Roles of AI Agents in Human-AI Co-creativity
@@
Desired roles of AI in human-AI collaboration have been widely discussed (maher2012computational, ; lubart2005can, ; negrete2014apprentice, ; li2023ai, ; li2023far, ) from different perspectives (e.g. categories of human-computer interaction for creativity (lubart2005can, ), tasks in ideation (maher2012computational, ), computational creativity (negrete2014apprentice, ), work division in data storytelling (li2023ai, ; li2023far, )). While Lubart et al. (lubart2005can, ) stressed the diversity of users for AI roles, the characteristics of target users are under-discovered. Through categorizing users (Sec.  5.4.2), We propose four distinct roles for AI agents. We mainly consider dimensions, namely creativity and accuracy, because both are the primary attributes that participants value in AI-assisted graphic design (Fig. 11). Creativity refers to generating unconventional design ideas, while accuracy pertains to fulfilling design expectations. Figure 11 illustrates the estimated levels of creativity and accuracy for each of the four AI roles, along with a guideline that deems both creativity and accuracy to be equally important. When AI agents serve as experts, designers with ill-defined tasks expect the highest level of creativity and demonstrate a relevant tolerance toward accuracy, in search of professional suggestions. However, participants’ expectations regarding the creativity of AI colleagues vary, but on average, they demand less creativity compared to AI experts. When it comes to AI colleagues, users prioritize diverse perspectives over novelty. Accuracy is also deemed equally important as effective communication between users and AI colleagues is crucial. Conversely, AI optimizer and AI executor roles demand less creativity as users prioritize satisfaction over surprise. The AI optimizer role allows for a moderate level of accuracy tolerance, falling between the levels expected from AI expert and AI colleague, since it does not involve input requirements. The role of executor necessitates the highest level of accuracy in executing users’ commands, leaving little room for creativity. In a human-AI co-creativity system, these AI roles may need to be implemented simultaneously or shifted due to the diverse needs of users. There is no one-size-fits-all user, but rather a range of user profiles that can be further categorized for clarity and customization purposes. Given the diversity and consistency of design challenges, users’ intentions are multifaceted and may evolve throughout the creative design process. Consequently, the development of a specific system must be meticulously planned in alignment with the intended user base and creative design scenarios. For example, Study II highlights a divergent perspective towards OptiMuse and Copilot, indicating that the roles of AI should be accompanied by a redesign of user interfaces. The requirements for AI in communication and selective choice should be adjusted accordingly.
@@
Figure 11. Participants’ expectations regarding the levels of creativity and accuracy required for AI agents’ roles. Creativity: Expert ¿ Colleague ¿ Optimizer ¿ Executor; Accuracy: Executor ¿ Colleague ¿ Optimizer ¿ Expert.
@@ 6.3. Limitations and Future Work
One limitation of our research is the scope of our user studies, which focused solely on graphic design as a creative design scenario and recruited participants with expertise in this field. This restricted user sample may have resulted in incomplete findings. Therefore, we view our guidelines for the nonlinear human-AI co-design framework as a starting point to enhance human-AI collaboration by emulating human collaborations. We aim to inspire future research in various domains, such as creative writing, music composition, drawing, coding, and design ideation.
@@
Admittedly, the importance of natural multimodal communication in human-human collaboration cannot be ignored. When it comes to communication methods, both research (pan_human-computer_2023, ; shi2023Understanding, ), design practices, and our formative studies have observed designers employing various forms, including text, voice, gestures, sketches, graphical operations, and even eye movements, to convey and exchange design ideas and reflections. We acknowledge that the time-consuming nature of text-based interaction hindered the efficiency and user experience of human-AI collaboration, making it less comparable to human-human collaboration. However, as highlighted by Ashktorab et al. (Ashktorab2020human, ), social perceptions exhibit significant differences when considering interaction with an AI compared to interacting with a human (Ashktorab2020human, ). To summarize design requirements without falling into stereotypes of AI agents, we determine to gain insights directly from human-human collaboration; and to make it comparable to the existing AI tools, we chose text-based communication as the primary mode in our proof-of-concept prototype. Shi et al. (shi2023Understanding, ) suggest that multimodal interaction in human-AI co-design is still in its early stages. Looking ahead, we anticipate that ongoing research on multimodal interaction will lead to more flexible communication with AI systems.
@@ 7. Conclusion
This research focuses on the collaboration between humans and AI in creative design. Based on a formative study, we have established design requirements and guidelines that can inspire the development of nonlinear AI-assisted tools. These guidelines aim to incorporate non-linear characteristics into human-AI collaboration and address the limitations of human-human collaboration. Following these guidelines, we propose a novel human-AI co-design framework and develop a proof-of-concept prototype called OptiMuse. To evaluate our framework, we conduct a comparative study and implement a baseline tool that follows a linear human-AI workflow. The findings demonstrate that OptiMuse significantly improves users’ task completion rates and provides valuable insights for future AI-powered tools.
@@ Acknowledgements.
Appendix A Appendix: Participants Feedback in Study I
A.0.1. The Inherent Nonlinearity of Requirements.
  • Requirement Ambiguousness Similarly, certain participants expressed their grievances drawing upon previous experiences: “When a client doesn’t have any design know-how, it’s hard for them to put their finger on what exactly the problem is. All they can really do is share their gut feelings about it.” -B1
@@ Appendix A Appendix: Participants Feedback in Study I
A.0.2. Pros and Cons for the Nonlinear Human-human Co-design Process.
@@
    “When I work alone, I tend to get stuck in my own ways and miss out on learning new methods of presenting information. Take the ‘timeline’ format in the experiment as an instance. While it may seem straightforward, I never really considered placing the three phases on a timeline. Collaborating with others opened my eyes to this ‘innovative’ approach, which served as a valuable knowledge enhancement.”-B4
@@
    “It reduces tangling. If I do it alone, sometimes I’m always struggling whether it’s going to work or not. But if there’s another person to discuss it with, he can guide me, like, ‘Let’s line up the worst piece first, and then we’ll go back to struggling with the details of it.’ ” - A4
    Misunderstandings are exposed and fixed swiftly through Q&A. “The communication process can help you to clarify your own modification needs, which are clearer and may be subject to change” - B6
    “It might require more than a couple of questions to really get what I’m trying to say, you know? It takes a bit of digging to grasp my point.” - A6 “Small misunderstandings can sometimes inspire new ideas for modifications. For example, just now when he said one color for this module and another color for the other. I interpreted it as one color for each row because I hadn’t thought about color matching module by module. But in fact, both of these modification options are quite good.” - B6
@@
Appendix B Appendix: Instructions for the Wizard
@@
Table 5. Standard Operations for OptiMuse and Copilot
@@
Actions. Chat in Figma; Exporting images from PowerPoint, Replacing images in Figma; Global Image changes in Figma; PowerPoint editing
@@
    Address the user as “Ning” (formal).
@@
    Use your judgment and offer proactive modification recommendations.
@@
    When the user changes their requirements, remind them of the changes. Op: “Previously, you requested a darker background color. Now, you’re asking for a warmer background color. Would you like to make the background color warmer while also darkening it?

### 20 Mar 2024 v3

@@ 5.4. Lessons We Learnt from Study II
What altered people’s preferences towards OptiMuse and Copilot? To answer this problem, we summarized people’s different expectations of AI agents in human-AI co-design and their preferences. Specifically, we try to address two questions inspired by Li et al. (Li et al., 2023b): (i)What tasks are anticipated from AI agents, and (ii)What roles are expected of AI agents?
@@ 5.4.1. What tasks are anticipated from AI agents?
We adjusted the sequence of these two questions because tasks will be frequently discussed in formulating the definition of AI agents’ roles. The tasks are as follows: Requirement Understanding: comprehend users’ design requirements conveyed through textual input; Requirement Analysis: engage in a dialogue with users to establish consensus regarding design objectives or operations. In other words, provide textual feedback towards users’ textual input; Solution Exploration: present alternative design options in either textual or visual formats; Review: evaluate designs based on design knowledge, making comparisons with existing solutions; Execution: transform operational requirements into visual outputs.
@@ 5.4.2. What roles are expected of AI agents?
Based on the interviews, our users expressed varying preferences and requirements regarding OptiMuse and Copilot. Specifically, they held different perceptions of the roles AI should play in graphic design. We categorized these roles as expert, colleague, optimizer, and executor, noting a diminished emphasis on creativity.
@@
  • Executor: Do just as I tell you. Among the 3/12 participants, the prevalent choice was for AI to work only as an executor. An executor AI is expected to faithfully carry out users’ commands. This specific type of user has clear goals and the possible operations required to achieve them, although they remain uncertain about the outcome until they actually put a hand to it. As a result, they depend on AI to reduce the manual effort involved in turning operations into visible results. They prefer to avoid requirement analysis, finding it cumbersome to explain operations to AI. Furthermore, they have no inclination to make selective choices because they wish to avoid the additional time and effort required for comparisons and selections among choices.
The tasks and adaptations within human-AI co-creativity systems corresponding to the expected roles of AI agents are listed in Table  3 and Table  4.
@@ 6. Discussion
6.1. Nonlinear Design for Co-creative Systems
@@
In popular tools like Copilot (Microsoft, 2023) and Midjourney (Midjourney, 2022), AI agents necessitate precise instructions and do not facilitate the remixing of various options (Fig. 3). They prioritize addressing the constraints inherent in the human-human co-design process (DR4, DR5). As described in Kulkarni et al.’s work (Kulkarni et al., 2023), designers often refine their prompts multiple times and regenerate images until they achieve relatively satisfactory results. Following this, they may choose to adopt either a portion or the entire outcome, either through manual editing or by blending prompts to generate new designs. Such turn-based linear workflow hinders the natural nonlinear navigation among creative design stages (Hatchuel and Weil, 2003; O’Connor, 1991; Council, 2024; Gero, 1990; Gero and Kannengiesser, 2004; Howard et al., 2008), which involves drawing inspiration from various parallel iterations and potentially modifying initial ideas along the way (Girotto, 2016). As a result, existing AI-assisted tools fail to consider the design requirements (Table.  1) because of their requirements for precise commands (fails to fulfill DR1-2) and restricted operations on outcomes (DR3).
@@
Existing frameworks (NEGRETE YANKELEVICH and MORALES ZARAGOZA, 2014; Kantosalo and Toivonen, 2016; Guzdial and Riedl, 2019) on human-AI collaboration have shifted attention from AI as mere tools to more collaborative and creative agents. However, they simply describe the co-design process as “turns” or “iteration”. We bridge this gap by formulating a nonlinear human-AI co-design framework featuring flexible communication and selective choices. Following the nonlinear framework, OptiMuse enables parallel iteration (Girotto, 2016) and remixing (Tan et al., 2022; Cheliotis et al., 2014; Oehlberg et al., 2015; Yue et al., 2019). Similar to Rezwana’s finding in Penpal (Rezwana and Maher, 2022), we observe enhanced user engagement in OptiMuse, where AI was perceived as more collaborative than a mere tool. In summary, our framework serves as a compelling example of integrating theories (creative design process, human-AI framework) and practices (collective creativity strategies, Study I), which offers practical guidelines for the development of future human-AI co-creative systems.
6.2. Roles of AI Agents in Co-creativity
Desired roles of AI in human-AI collaboration have been widely discussed from different perspectives (e.g. categories of human-computer interaction for creativity (Lubart, 2005), tasks in ideation (Maher, 2012), computational creativity (NEGRETE YANKELEVICH and MORALES ZARAGOZA, 2014), work division in data storytelling (Li et al., 2023b, a)). While Lubart et al. (Lubart, 2005) stressed the diversity of users for AI roles, the characteristics of target users are under-discovered.
@@ 6.2. Roles of AI Agents in Co-creativity
Categorizing users’ expected AI collaborators (Sec.  5.4.2), we propose four distinct roles for AI agents. We mainly consider dimensions, namely creativity and accuracy, because both are the primary attributes that participants value in AI-assisted graphic design (Fig. 9). Creativity refers to generating unconventional design ideas, while accuracy pertains to fulfilling design expectations. Fig. 9 illustrates the estimated levels of creativity and accuracy for each of the four AI roles, along with a guideline that deems both creativity and accuracy to be equally important. When AI agents serve as experts, designers with ill-defined tasks expect the highest level of creativity and demonstrate a relevant tolerance toward accuracy, in search of professional suggestions. However, participants’ expectations regarding the creativity of AI colleagues vary, but on average, they demand less creativity compared to AI experts. When it comes to AI colleagues, users prioritize diverse perspectives over novelty. Accuracy is also deemed equally important as effective communication between users and AI colleagues is crucial. Conversely, AI optimizers and executors demand less creativity as users prioritize satisfaction over surprise. The AI optimizer allows for a moderate level of accuracy tolerance, falling between the levels expected from AI experts and AI colleagues, since it does not involve input requirements. The role of the executor necessitates the highest level of accuracy in executing users’ commands, leaving little room for creativity. In a human-AI co-creativity system, these AI roles may need to be implemented simultaneously or shifted due to the diverse needs of users. There is no one-size-fits-all user, but rather a range of user profiles that can be further categorized for clarity and customization purposes. Given the diversity and consistency of design challenges, users’ intentions are multifaceted and may evolve throughout the creative design process. Consequently, the development of a specific system must be meticulously planned in alignment with the intended user base and creative design scenarios. For example, Study II highlights a divergent perspective towards OptiMuse and Copilot, indicating that the roles of AI should be accompanied by a redesign of user interfaces and AI actions in communication and selective choice should be adjusted accordingly.
@@
Figure 9. Participants’ expectations regarding the creativity and accuracy required for different roles of AI agents.
@@ 6.3. Limitations and Future Work
One limitation of our research is the scope of user studies, which focused solely on graphic design as a creative design scenario and recruited participants with expertise in this field. This restricted user sample may have resulted in incomplete findings. Therefore, we view our guidelines for the nonlinear human-AI co-design framework as a starting point to enhance human-AI collaboration by emulating human collaborations. We aim to inspire future research in various domains, such as creative writing, music composition, drawing, coding, and design ideation.
@@
Admittedly, the importance of natural multimodal communication in human-human collaboration cannot be ignored. When it comes to communication methods, both research (Pan et al., 2023; Shi et al., 2023), design practices, and our formative studies have observed designers employing various forms, including text, voice, gestures, sketches, graphical operations, and even eye movements, to convey and exchange design ideas and reflections. We acknowledge that the time-consuming nature of text-based interaction hindered the efficiency and user experience of human-AI collaboration, making it less comparable to human-human collaboration. However, as highlighted by Ashktorab et al. (Ashktorab et al., 2020), social perceptions exhibit significant differences when considering interaction with an AI compared to interacting with a human (Ashktorab et al., 2020). To summarize design requirements without falling into stereotypes of AI agents, we determined to gain insights directly from human-human collaboration; and to make it comparable to the existing AI tools, we chose text-based communication as the primary mode in our POC. Shi et al. (Shi et al., 2023) suggest that multimodal interaction in human-AI co-design is still in its early stages. Looking ahead, we anticipate that ongoing research on multimodal interaction will lead to more flexible communication with AI systems.
@@ 7. Conclusion
This research focuses on the collaboration between humans and AI in creative design. Based on a formative study, we have established design requirements and guidelines that can inspire the development of nonlinear AI-assisted tools. These guidelines aim to incorporate nonlinear characteristics into human-AI collaboration and address the limitations of human-human collaboration. Following these guidelines, we propose a novel human-AI co-design framework and develop a POC, OptiMuse. To evaluate our framework, we conduct a comparative study and implement a baseline tool that follows a linear human-AI workflow. The findings demonstrate that OptiMuse significantly improves users’ task completion rates and provides valuable insights for future AI-powered tools.
@@ Acknowledgements.
The work was supported by NSF of China (U22A2032, 62302440) and the Collaborative Innovation Center of Artificial Intelligence by MOE and Zhejiang Provincial Government.
Appendix A Participants Feedback in Study I
A.1. The Inherent Nonlinearity of Requirements.
  • Requirement Ambiguity Participants expressed their grievances drawing upon previous experiences: “When a client doesn’t have any design know-how, it’s hard for them to pinpoint the problem. All they can do is share their gut feelings about it.” -B1
@@ Appendix A Participants Feedback in Study I
A.2. Pros and Cons for the Nonlinear Human-human Co-design Process.
@@
    “When I work alone, I tend to get stuck in my ways and miss out on learning new methods of presenting information. Take the ‘timeline’ format in the experiment as an instance. While it may seem straightforward, I never considered placing the three phases on a timeline. Collaborating with others opened my eyes to this ‘innovative’ approach, which served as a valuable knowledge enhancement.”-B4
@@
    “It reduces tangling. When designing alone, I’m always struggling with whether it’s going to work or not. But if there’s another person to discuss, he can guide me, like, ‘Let’s line up the worst piece first, and then we’ll go back to struggling with the details of it.’ ” - A4
    Misunderstandings are exposed and fixed swiftly through Q&A. “The communication process helps to clarify your own modification needs, which are clearer and may be subject to change” - B6
    “It might require more than a couple of questions to get what I’m trying to say.” - A6
    “Misunderstandings sometimes inspire new ideas for modifications. For example, just now when he said one color for this module and another color for the other. I interpreted it as one color for each row because I hadn’t thought about color matching module by module. But in fact, both of these modification options are quite good.” - B6
@@
Table 5. Standard Operations for OptiMuse and Copilot
@@
Appendix B Instructions for the Wizard
@@
Actions. Chat in Figma; Export images from PowerPoint; Replace images in Figma; Change the Preview in Figma; Edit PowerPoint
@@
    Address the user as “Nin” (formal).
@@
    Use your own judgment and offer proactive modification recommendations.
@@
    When the user changes their requirements, remind them of the changes. Op: “Previously, you requested a darker background color. Now, you’re asking for a warmer background color. Would you like to make the background color warmer and darker?